### v1

# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

Abstract
The recent discovery of superconductivity with a transition temperature $T_{c}$ over 40 K in La3Ni2O7 and (La,Pr)3Ni2O7 thin films at ambient pressure marks an important step in the field of nickelate superconductors. Here, we perform a renormalized mean-field theory study of the superconductivity in $\mathrm{La_{3}Ni_{2}O_{7}}$ thin films, using a bilayer two-orbital $t-J$ model. Our result reveals an $s_{\pm}$-wave pairing symmetry driven by the strong interlayer superexchange coupling of $d_{z^{2}}$ orbital, resembling the pressurized bulk case. Also, we roughly reproduce the experimentally reported nodeless shape of the superconducting gap at the $\beta$ pocket and the superconducting $T_{c}$. To gain insight into the orbital feature of superconductivity, we explore the projection of different pairing bonds on Fermi surface. We find that the nodeless gap at $\beta$ pocket is related to the interlayer pairing within both $d_{z^{2}}$ and $d_{x^{2}-y^{2}}$ orbitals, in which the latter is triggered by the former through hybridization, and both hold the same sign.
preprint: APS/123-QED
Introduction.--The discovery of superconductivity with a transition temperature $T_{c}$ near 80 K in the Ruddlesden-Popper (RP) bilayer nickelate La3Ni2O7 (LNO)  [1] has gained widespread interests in the field of high-$T_{c}$ superconductivity, and it was soon followed by another observation of superconductivity in trilayer nickelate La4Ni3O10 with $T_{c}$=20-30 K [2, 3], both under high pressure. The discovery of RP nickelate-based superconductors has marked a breakthrough in the field of high-$T_{c}$ superconductivity, and therefore raised considerable subsequent investigations both theoretically [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38] and experimentally [39, 40, 41, 42, 43, 44, 45, 46]. However, the need for high pressure has largely confined some critical experimental techniques for a direct probe of the superconducting phase, such as angle resolved photoemission spectroscopy (ARPES) and scanning tunneling microscopy (STM), which prevents a decisive confirmation of the superconducting pairing among various theoretical proposals [16, 34]. On the other hand, achieving ambient-pressure superconductivity is a crucial step for practical considerations, and in this respect, there are some suggestions on the possible route, such as using chemical substitute [47, 48, 49] or applying strain [50, 51, 52, 53].
Very recently, superconductivity with $T_{c}$ over 40 K in LNO [54] and (La,Pr)3Ni2O7 [55] thin films at ambient pressure have been reported. The X-ray absorption spectroscopy (XAS) analysis and scanning transmission electron microscopy (STEM) reveal that, the Ni ions in LNO thin films maintain a mixed valence state and the apical Ni-O-Ni bond angle approaches 180 [54], which closely resembles the high-pressure bulk phase of LNO. The ARPES measurement [56] in thim films also confirms the dominant Ni-$d_{x^{2}-y^{2}}$ and $d_{z^{2}}$ orbital spectral weights around the Fermi energy, which is consistent with DFT calculations [57, 58, 59, 53, 52]. Moreover, a direct ARPES measurement of the superconducting electronic structure on the Fermi surface (FS) was performed recently [60], which reveals a notable gap opening on the $\beta$ FS sheet and without showing any node along the Brillouin zone (BZ) diagonal direction. This observation is accompanied by another STM measurement on superconducting state [61], which shows a two-gap structure on the FS and a fitting to such gaps demonstrates a preferable anisotropic $s$-wave pairing. The two results are generally prone to the idea of the predominant $s_{\pm}$-wave pairing symmetry in the system as predicted by several theoretical studies [18, 16, 7, 35, 58].
Figure 1: (a) Band structure of the 4-orbital tight-binding model adopted from the DFT simulation of the double-stacked $\mathrm{La_{3}Ni_{2}O_{7}}$ thin film [57]. The red dashed line denotes the undoped case from DFT while the bule one denotes the slightly electron doped case as is used in our study. (b) Schematic of $d_{x^{2}-y^{2}}$ (red) and $d_{z^{2}}$ (blue) orbitals of the model. Two major superexchange couplings $J_{\bot}$, $J_{||}$ are drawn. See also the main text.
In this paper, using a bilayer $t-J$ model that captures the key ingredient of LNO thin films, we perform a renomalized mean-field theory (RMFT) [7] calculation of the superconductivity. Our result reveals an $s_{\pm}$-wave pairing and reproduces the nodeless shape of the superconducting gap at $\beta$ pocket as reported in Ref. [60]. Apparently, the presence of a finite gap along BZ diagonal direction has indicated a direct $s_{\pm}$-wave pairing on $d_{x^{2}-y^{2}}$ orbital. In general, the gap structure on Fermi surface is sensitive to the relative strengths of pairing configurations in real space, and our result suggests that such a gap shape is a joint effort of the interlayer pairing bonds within both $d_{z^{2}}$ and $d_{x^{2}-y^{2}}$ orbitals.
Model and method.--The bilayer two-orbital $t-J$ model is defined as
$\mathcal{H}=$ $\mathcal{H}_{t}+\mathcal{H}_{J},$ (1)
$\mathcal{H}_{t}=$ $\sum_{ij}\sum_{\mu\nu}\sum_{\sigma}t_{ij}^{\mu\nu}c_{i\mu\sigma}^{\dagger}c_{j\nu\sigma}-\mu\sum_{i\mu\sigma}n_{i\mu\sigma}$ (2)
$\mathcal{H}_{J}=$ $J_{\bot}\sum_{i}\boldsymbol{S}_{iz_{1}}\cdot\boldsymbol{S}_{iz_{2}}+J_{||}\sum_{\langle ij\rangle}^{\mu=x_{1},x_{2}}\boldsymbol{S}_{i\mu}\cdot\boldsymbol{S}_{j\mu}$ (3)
$+$ $J_{xz}\sum_{\langle ij\rangle}^{\mu\nu=x_{1}z_{1},x_{2}z_{2}}\boldsymbol{S}_{i\mu}\cdot\boldsymbol{S}_{j\nu}-J_{H}\sum_{i}^{\mu\nu=x_{1}z_{1},x_{2}z_{2}}\boldsymbol{S}_{i\mu}\cdot\boldsymbol{S}_{i\nu}.$
In this model, the basis is $\Psi=(d_{z_{1}},d_{x_{1}},d_{z_{2}},d_{x_{2}})^{T}$ that contains $d_{z^{2}}$ and $d_{x^{2}-y^{2}}$ orbitals for the first and second layers, as shown in Fig. 1(b). $\mathcal{H}_{t}$ is the tight-binding Hamiltonian adopted from DFT simulation of the double-stacked LNO thin films [57]. Here, $ij,\mu\nu,\sigma$ denote the lattice, orbital, and spin indices, respectively. The band structure is shown in Fig. 1(a), in which a small electron doping level with chemical potential $\mu=-0.05$ is used in our calculation to better agree with the experiments, which gives the elctron filling $n_{z}$=0.8, $n_{x}$=0.58. For the Heisenberg term $\mathcal{H}_{J}$, $J_{\bot}$ is the interlayer superexchange coupling of $d_{z^{2}}$ orbitals, and $J_{||},J_{xz}$ is the inplane nearest neighbor superexchange coupling of $d_{x^{2}-y^{2}}$ orbital, and superexchange coupling between $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$, respectively. Also, the Hund’s coupling $J_{H}$ is included for generality. To quantitatively estimate $J_{\bot},J_{||},J_{xz}$, we perform the exact diagonalization (ED) calculation. Taking $J_{\bot}$ as an example, ED is performed for a 5-site chain containing two $d_{z^{2}}$ and three $p_{z}$ orbitals along the $z$ axis, where the $p-d$ hoppings are also obtained from the DFT calculation [57]. Under the atomic subspace ($N$=8, $S_{z}$=0) that corresponds to half-filling of $d_{z^{2}}$ orbital, $J_{\bot}$ is determined as the energy difference of the lowest singlet and triplet states [8]. We finally obtain $J_{\bot}$=0.135 eV, $J_{||}$=0.084 eV, $J_{xz}$=0.03 eV. Apart from this, a realistic $J_{H}$=1 eV is also used in this study.
The above $t-J$ model is solved under the RMFT. The RMFT [62] follows a conventional mean-field decomposition of spin exchange, which yields the order parameters $\chi_{ij}^{\mu\nu}=\langle c^{\dagger}_{i\mu\uparrow}c_{j\nu\uparrow}\rangle$ and $\Delta^{\mu\nu}_{ij}=\langle c^{\dagger}_{i\mu\uparrow}c^{\dagger}_{j\nu\downarrow}\rangle$ under the spin symmetry. Besides, $\mathcal{H}_{t}$ and $\mathcal{H}_{J}$ are respectively coupled by the Gutzwiller renormalization factors $g_{t}$ and $g_{J}$ to recover the strong renormalization from correlation [63]. The obtained quadratic Hamiltonian is then claculated self-consistently, during which the electron densities are fixed to certain level. The eventual pairing symmetry can be determined from the phase structure of various pairing bonds $\Delta^{\mu\nu}_{ij}$. We refer the reader to Ref. [7] for a comprehensive introduction of the method.
Figure 2: (a) Superconducting gap $g_{t}\Delta$ projected onto FS for $\mathrm{La_{3}Ni_{2}O_{7}}$ thin films. (b-d) $g_{t}|\Delta|$ as a function of angle $\theta$ for $\alpha$, $\beta$ and $\gamma$ pockets, respectively. $\theta$ is indicated in (a).
Pairing symmetry and superconducting gap.-- Figure. 2(a) presents the superconducting gap matrix $[g_{t}\Delta]_{\mu\nu}$ projected onto FS, which demonstrates that $\beta,\gamma$ pockets are nodeless and of opposite sign, while $\alpha$ pocket shows a sign flip. To better illusstrate the gap features, Fig. 2(b-d) show the absolute value $g_{t}|\Delta|$ as a function of angle $\theta$ for $\alpha,\beta$ and $\gamma$ pockets, respectively. For $\alpha$ pocket in Fig. 2(b), the gap demonstrates a node between $\theta=45^{\circ}\sim 90^{\circ}$. For $\beta$ pocket in Fig. 2(c), the gap shows a parabolic shape centering at $\theta=45^{\circ}$, which gets slightly decrease when approaching the corner. Note that the gap shape as well as its magnitude around the BZ diagonal direction from our result are generally coordinated with the ARPES result [60], which suggests that our RMFT captures the key ingredient in the LNO thin films. For $\gamma$ pocket in Fig. 2(d), we find the gap magnitude is around 5 times larger than that in Fig. 2(b), and a minimum is also identified at the BZ diagonal direction. Apparently, such large gap opening at $\gamma$ pocket should be attributed to the strong interlayer superexchange coupling $J_{\bot}$, which is also consistent with previous calculations for bulk LNO [18, 16, 25, 7]. Overall, the pairing structure in Fig. 2 should be described as the $s_{\pm}$-wave, which can also be seen from our result that besides the formation of the large interlayer pairng bond for $d_{z^{2}}$ orbital ($\Delta_{\bot}^{z}$), the inplane pairing bonds have the same magnitude when along $x,y$ directions for both $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$ orbitals ($\Delta_{||}^{x},\Delta_{||}^{z})$. Note that, in our $t-J$ model only the pairing bonds $\Delta_{\bot}^{z},\Delta_{||}^{x},\Delta_{||}^{xz}$ are associated with the applied $J_{\bot},J_{||},J_{xz}$, respectively, while other terms like the interlayer pairing of $d_{x^{2}-y^{2}}$ orbital ($\Delta_{\bot}^{x}$) do not have the corresponding $J$ and should not appear in the calculation. However, they can still gain finite values from the mean-field wave function as $\Delta_{ij}^{\mu\nu}=\langle n{\rm k}|c_{i\mu\uparrow}^{\dagger}c^{\dagger}_{j\nu\downarrow}|n{\rm k}\rangle$, which reflects the pairing tendency under a global coherent state. For this reason, in Fig. 2 we have incorporated all the relevant $\Delta$, and in the following we will illustrate the role of different $\Delta$ on FS.
Orbital-resolved superconducting gap.-- Figure. 3 presents the orbital-resolved superconducting gap for (a) $\Delta_{\bot}^{z}$, (b) $\Delta_{||}^{z}$, (c) $\Delta_{\bot}^{x}$, and (d) $\Delta_{||}^{x}$, respectively. For $\Delta_{\bot}^{z}$ in Fig. 3(a), the gap is positive for the bonding branch of $\alpha,\gamma$ pockets, and is negative for the anti-bonding branch of $\beta$ pocket. For $\Delta_{||}^{z}$ in Fig. 3(b), the gap structure clearly demonstrates the sign-flip $s_{\pm}$-wave with the node appearing at the reduced BZ boundary. Also, another node at the BZ diagonal direction in both Fig. 3(a,b) is identified due to the vanishing of hybridization. For $\Delta_{\bot}^{x},\Delta_{||}^{x}$ in Fig. 3(c,d), similar behavior is observed as compared to Fig. 3(a,b), but with a much smaller magnitude. Notably, the sign in Fig. 3(d) is opposite to that in Fig. 3(b), namely, $sgn(\Delta_{||}^{x})=-sgn(\Delta_{||}^{z})$. This feature is also seen in a slave-boson mean-field study [31], and can be explained as a result of the competition between $J_{\bot}$ preferred $s_{\pm}$-wave and $J_{||}$ preferred $d$-wave pairings [64]. Finally, combining Fig. 3(a,c), it is evident that the experimentally observed nodeless gap at $\beta$ pocket is a joint effort of the interlayer pairing within both $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$ orbitals. More insights on gap structures with the real space pairing configurations are provided in supplementary [64].
Transition temperature.--In Fig. 4, we show the superconducting gap as a function of temperature $T$, which reveals two dominated pairing bonds from $d_{z^{2}}$ orbitals (red), in line with Fig. 3. As increasing $T$, these gaps decrease in a mean-field manner and simultaneously drop to zero at $T_{c}$$\approx$60 K. Such value is comparable to the experimental reported $T_{c}$=40$\sim$60 K in the thin films [54, 55, 50]. Given that our previous RMFT study also obtains an experimental comparable $T_{c}\approx 80$ K for the pressurized bulk LNO [7], it is reasonable to believe that the decrease of $T_{c}$ in the thin films should originate from the decrease of $J_{\bot}$. This speculation draws us to the difference in the electronic structure, especially in the mainly concerned interlayer hopping $t_{\bot}^{z}$ for $d_{z^{2}}$ orbital. Intriguingly, we find that the $J_{\bot}$=0.135 eV in this study is $\sim 25\%$ smaller than that of $J_{\bot}$=0.18 eV for the pressurized bulk [8] as estimated from ED, which is coordinated with a simple estimation using $J_{\bot}\propto(t_{\bot}^{z})^{2}$ [4, 57]. Recently, a strain-tuning experiment in LNO thin films has reported an increase of $T_{c}$ from 10 K in the tensile-strained $\mathrm{SrTiO_{3}}$ substrate to 60 K in the compressively strained $\mathrm{LaAlO_{3}}$ substrate under 20 GPa of pressure [50]. This indicates a positive dependence of $T_{c}$ on the lattice ratio $c/a$ and, at first glance, violates the above understanding. However, we believe $J_{\bot}$ can only determine the upper bound of $T_{c}$, while other factors like the hybridization, density wave, and interlayer Josephson coupling can also have intricate impact, as also pointed out in Ref. [65, 66, 67, 68, 69, 50, 70].
Figure 3: Orbital-resolved superconducting energy gap on FS for (a) $\Delta_{\bot}^{z}$, (b) $\Delta_{||}^{z}$, (c) $\Delta_{\bot}^{x}$, and (d) $\Delta_{||}^{x}$. $g_{t}^{z/x}$ is the renormalization factor for $d_{z^{2}}/d_{x^{2}-y^{2}}$ orbital. $\Delta_{\bot/||}^{z/x}$ denotes the interlayer/inplane pairing bond for $d_{z^{2}}/d_{x^{2}-y^{2}}$ orbital.
Figure 4: Superconducting gap as a function of temperature for different pairing bonds. $g_{t}^{z/x}$ is the renormalization factor for $d_{z^{2}}/d_{x^{2}-y^{2}}$ orbital, and $\Delta_{\bot/||}^{z/x}$ is the interlayer/inplane pairing bond for $d_{z^{2}}/d_{x^{2}-y^{2}}$ orbital.
Discussion.-- Our RMFT study of the superconductivity in LNO thin films has predicted an $s_{\pm}$-wave pairing symmetry. The result largely resembles that of the pressurized bulk LNO [18, 16, 25, 7] giving the similarity in the electronic structure. Mathematically, our application of Gutzwiller projection  [62, 18] on a multi-orbital $t-J$ model emposes a constraint on the renormalization condition so as to bring a simplification to have the quadratic form of the Hamiltonian, hence can be solved self-consistently like the standard Bardeen-Cooper-Schrieffer (BCS) theory. For this reason, it is expected that the obtained $s_{\pm}$-wave mainly reflects pairing preference for the FS toplogy in LNO [71]. Nevertheless, we find that the $s_{\pm}$-wave is quite robust under various approaches from weak to strong couplings, which convinces us that the fluctuation of correlated atomic states might not significantly alert such picture. In this picture, the $\gamma$ pocket plays a central role. Experimentally, there are two relevant ARPES measurements in which the $\gamma$ pocket shows visible profile for that performed on the normal state [56] but invisible for that on the superconducting state [51]. From our result, this contradiction might be due to the opening of a large gap that exceeds the detected energy range, as shown Fig. 2(d). Further experimental evidences are expected on this issue.
Conclusion-- To summarize, we perform a RMFT study of the superconductivity in the LNO thin films using a bilayer $t-J$ model. Our result reveals an $s_{\pm}$-wave pairing symmetry driven by the interlayer superexchange coupling, resembling the bulk LNO case. Our result also roughly reproduces the experimentally reported nodeless shape of the superconducting gap at $\beta$ pocket  [60] as well as the superconducting $T_{c}$. In addition, we systematically analyse the relation between orbital-resolved pairing bonds and the gap structure on FS. The result shows that the nodeless gap of $\beta$ pocket is related to the interlayer pairing within both $d_{z^{2}}$ and $d_{x^{2}-y^{2}}$ orbitals, in which the latter is triggered by the former through hybridization. Both of them hold the same sign.
We thank the useful discussion with Wéi Wú. This project was supported by NSFC-12494591, NSFC-92165204, NKRDPC-2022YFA1402802, Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices (2022B1212010008), Guangdong Fundamental Research Center for Magnetoelectric Physics (2024B0303390001), and Guangdong Provincial Quantum Science Strategic Initiative (GDZX2401010).